OpenStudy (anonymous):

The length of a rectangle is 3 cm more than 10 times its width. The area of the rectangle is 75 cm2. A quadratic function can be used to represent the dimensions of the rectangle. What quantity of the rectangle represents the domain of that function?

OpenStudy (anonymous):

let its width be x than length= 3 + 10x breadth= 1x then area=> 3+10x * x =75

OpenStudy (anonymous):

3x + 10xsq - 75 = 0

OpenStudy (anonymous):

then use quadratic formula

OpenStudy (anonymous):

domain of the function is width